OpenStudy (anonymous):

easy question please help.. What is the slope between the two points (4, 13) and (1, 4)? 3 -1/3 1/3 -3

OpenStudy (anonymous):

take the ratio of the (difference in y's) / (difference in x's)

OpenStudy (anonymous):

so its B

OpenStudy (anonymous):

(difference in y's) / (difference in x's) = (13-4)/(4-1)

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

are you guessing?

OpenStudy (anonymous):

educated guesss

OpenStudy (anonymous):

you don't need to guess... just calculate what I gave you... (difference in y's) / (difference in x's) = (13-4)/(4-1)

OpenStudy (anonymous):

more help please

OpenStudy (anonymous):

13-4= 4-1= do these...

OpenStudy (anonymous):

9,3

OpenStudy (anonymous):

so (13-4)/(4-1) = 9/3 =

OpenStudy (anonymous):

yeah then

OpenStudy (anonymous):

ok?...:)

OpenStudy (anonymous):

answers 3

OpenStudy (anonymous):

yes...:) see you didn't need to guess at this...:)

OpenStudy (anonymous):

can u help with this. What is the equation of the line that passes through the point (4, 1) and has a slope of 5? y = 5x - 3 y = 5x - 1 y = 5x + 3 y = 5x - 19

OpenStudy (anonymous):

ok.. \[\huge y=mx+b \] is the equation of a line in slope intercept-form.... we already have our slope (m), and (x,y)=(4,1). we need to find b.

OpenStudy (anonymous):

 click on the pencil icon in the upper right corner so you can fill in the boxes with the numbers that are given....